OpenStudy (anonymous):

A radiator contains 15 quarts of fluid, 30% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 35% antifreeze? (Round your answer to two decimal places.)

OpenStudy (anonymous):

\[\%Antifreeze = \frac{V_{Antifreeze}}{V_{Fluid}}\] \[V_{Fluid}=\frac{V_{Antifreeze}}{\%Antifreeze}\] \[V_{Antifreeze}=V_{fluid} * \%Antifreeze\] I want to find the amount of antifreeze in a 15 quart solution with 30% antifreeze \[V_{Antifreeze}=15*0.30\] =18/4 quarts of antifreeze Similarly, I want to find the amount of antifreeze in a 15 quart solution with 35% antifreeze first. \[V_{Antifreeze}=15*0.35\] = 21/4 quarts of antifreeze

OpenStudy (anonymous):

the difference between 21/4 and 18/4 is 3/4 quarts, which is the amount of pure antifreeze I've added in.

OpenStudy (anonymous):

SO the V_fluid I replaced with 3/4 quarts of antifreeze is (3/4)/ 0.35

OpenStudy (anonymous):

Therefore, I must replace 15/7 quarts of the fluid with pure antifreeze :-D

OpenStudy (anonymous):

But I'm mistaken

OpenStudy (anonymous):

It's telling me that I have the wrong answer.

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I think I figured it out now: \[\text{Antifreeze}\ \ \ \%100\ \ \ x\text{ quarts}\] \[\text{Fluid }\%30\ \ \ (15-x)\text{quarts}\] \[\text{New mixture}\ \ \ \ \%35\ \ \ \ 15quarts\]

OpenStudy (anonymous):

\[100x + 30(15-x)=35*15\] solve for x \[x = \frac{15}{14}\text{quarts}\]

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

therefore, 1.07 quarts of fluid must be drained and replaced with pure antifreeze

OpenStudy (anonymous):

Yeah, Thanks!

OpenStudy (anonymous):

Can you help me with my other word problem too?

OpenStudy (anonymous):

The one about investing.